OpenStudy (anonymous):

At the first meeting of the Table Tennis Club, the seven members decide to have a tournament in which each player will play a game against each other player. How many games will there be?

OpenStudy (anonymous):

i got 49 but i think thats wrong

OpenStudy (anonymous):

This is number of combinations of 7 things taken 2 at a time n!/r!(n-r)! = 7! / 2! (5!) = 7 * 6 / 2 = 21 Label players A to G A plays B to G: 6 games B plays C to G: 5 games C plays D to G: 4 games D plays E to G: 3 games E plays F to G 2 games F plays G: 1 game total: 21 games nice when it works out as it should!

OpenStudy (anonymous):

Good night!